OpenStudy (anonymous):

A system of equations is shown below: 6x – 5y = 5 3x + 5y = 4 The x-coordinate of the solution to this system of equations is _____.

OpenStudy (anonymous):

Solve equation 1 for y. Plug for y in equation 2. Solve for x.

OpenStudy (anonymous):

umm i dont know how to do that i just need help finishing im almost done with my first segment in flvs took me 20 something weeks to get here

OpenStudy (luffingsails):

Hi.

OpenStudy (luffingsails):

Can you try to solve equation 1 for y?

OpenStudy (anonymous):

Put the first equation in the form y= - _____x + number

OpenStudy (anonymous):

umm so lost can u give me the equation in the form first? that might help me

OpenStudy (anonymous):

Show me what you've done so far.

OpenStudy (anonymous):

well i've got y=6x+5 but im not sure if thats right

OpenStudy (anonymous):

well i've got y=6x+5 but im not sure if thats right

OpenStudy (luffingsails):

6x - 5y = 5 -5y = 5 - 6x

OpenStudy (luffingsails):

y = (5-6x)/(-5)

OpenStudy (luffingsails):

Now, take that equation and substitute it into the second equation. Then solve for x.

OpenStudy (luffingsails):

It's a bit of book work, but you can give it a try. I'll be back to help.

OpenStudy (anonymous):

uh wow o.o

OpenStudy (anonymous):

ok i got 6x-5=4-3x combine 6x and 3x and its 9x and add -5+4? or?

OpenStudy (anonymous):

You should get 9x - 5 = 4 , so you are correct. So 9x=9 , so x=?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem :)